GLOBAL IRREGULARITY FOR DEGENERATE ELLIPTIC OPERATORS

MICHAEL CHRIST

Abstract. Examples are given of degenerate elliptic operators on smooth, compact manifolds that are not globally regular in C^{∞}. These operators degenerate only in a rather mild fashion. Certain weak regularity results are proved, and an interpretation of global irregularity in terms of the associated heat semigroup is given.